A cube of wood supporting 200 gm mass just floats in water. When the mass is removed. The cube ruses by   
2 cm. What is the  size of the cube?

Grade:11

5 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
Let the edge of cube be ℓ. When mass is on the cube of wood
200 g + ℓ3 d wood g = ℓ3 dwater g
⇒ ℓ3 dwood = ℓ3 dwater – 200 … (i)
When the mass is removed
3 dwood = (ℓ - 2 ) ℓ2 dwater … (ii)
From (i) and (ii)
3 dwater - 200 = (ℓ- 2) ℓ2 dwater
But dwater = 1
∴ ℓ3 - 200 = ℓ2 (ℓ - 2)
⇒ ℓ = 10 cm
Thanks
Kevin Nash
askIITians Faculty
SHAIK AASIF AHAMED
9 years ago
Hello student,

Let the edge of the cube be x and the cube rises 2 cm on removing the mass. Thus, the weight of the body must be equal to the thrust provided by 2 cm height of the cube of base areas x2.

i.e. 200 g = (2 x2) 1 g

or x = 10 cm

Thanks and regards

Shaik Aasif

askIITians faculty

Gopivallabh
6 years ago

A CUBE OF wood floating in water supports 200g mass resting at centre of its toop face . when mass is removed, cube rises 2cm . find out volume of cube

 
Shreya
5 years ago
Let the edge of the cube be l cmWe know that mass=density× volumeVolume=area of the cube×heightGiven the cube rises by 2cm hence ht=2cmSo L×L×2×1g=200gm(density of water=1gm)L×L=100So l=10cm
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let the edge of the cube be x
cube rises 2 cm on removing the mass.
Thus, weight of the body must be equal to the thrust
provided by 2 cm height of the cube of base areas x2.
i.e. 200 g = (2 x2) 1 g
or x = 10 cm
Thanks and Regards